OpenStudy (nuccioreggie):

A set of equations is given below: Equation A: y = x + 1 Equation B: y = 4x + 5 Which of the following steps can be used to find the solution to the set of equations? x + 1 = 4x + 5 x = 4x + 5 x + 1 = 4x x + 5 = 4x + 1

OpenStudy (michele_laino):

hint: what is the step which has been written using the substitution method?

OpenStudy (mathmale):

that's a good hint. There are several different methods for solving systems of linear equations, including substitution, graphing, elimination by addition/subtraction, etc. Substitution would work well here, but so would addition/subtraction. Keeping this in mind, try applying the substitution method to the system at hand, and then try applying the elimination method to the same problem. By doing this you should be able to reject the answer possibilities that would not fit one of these methods or the other.

OpenStudy (nuccioreggie):

so its b

OpenStudy (michele_laino):

If I substitute the value of \(y\) from first equation, into the second one, I get: \[\huge x + 1 = y = 4x + 5\] so, what is the right option?

OpenStudy (michele_laino):

such expression is equivalent to this one: \[\huge x + 1 = 4x + 5\]

OpenStudy (nuccioreggie):

c

OpenStudy (michele_laino):

option C, contains a different equation

OpenStudy (nuccioreggie):

so d

OpenStudy (michele_laino):

you have to search for the option which contains this equation: \[\huge x + 1 = 4x + 5\]

OpenStudy (nuccioreggie):

A

OpenStudy (michele_laino):

that's right! Option A, it is the application of the substitution method